• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name W. Korsch

    Log entry time 19:36:03 on July19,2001

    Entry number 66778

    keyword=Target - EPR - 41.7%

    18:25 EPR measurements , important: see note below

    P(initial): 44.0 %
    P(final): 41.7% (average of last 2 flips)

    Conditions: beam off, target position: pick-up coil, lasers #5,6,7 on, grad. coils off,
    B-field: 0 deg.

    0th order fit:
    freq1: -48.11 kHz +/- 1.75 kHz | Polarisation: -43.93 % +/- 1.6 %
    freq2: 46.43 kHz +/- 1.58 kHz | Polarisation: 42.4 % +/- 1.45 %
    freq3: -45.66 kHz +/- 1.72 kHz | Polarisation: -41.7 % +/- 1.57 %
    freq4: 45.25 kHz +/- 1.78 kHz | Polarisation: 41.33 % +/- 1.63 %

    first order fit:
    freq1: -48.48 kHz +/- 1.68 kHz | Polarisation: -44.27 % +/- 1.53 %
    freq2: 45.73 kHz +/- 1.49 kHz | Polarisation: 41.76 % +/- 1.36 %
    freq3: -46.47 kHz +/- 1.76 kHz | Polarisation: -42.44 % +/- 1.61 %
    freq4: 45.12 kHz +/- 1.81 kHz | Polarisation: 41.21 % +/- 1.65 %

    Comments: While watching for spin flip to occur on target computer screen, noticed
    that no spin flip happened!!! Looked at oscilloscope and saw the RF was not
    displayed, looked at HP function generator: amplutide = 2.4 V(rms), green LED
    on. So that's ok. What was wrong!!! Toggled NMR/EPR switch on capacitor
    box again and then RF appeared!! (Switched from NMR to EPR position
    before EPR measurement). As soon as RF was on, spin flip happened and
    then RF went off -> ended up in wrong spin state!! Started immediately
    new sweep to flip spin back. Then I waited a little too long before I turned off
    the RF button, i.e. second spin flip happened. Performed another (1/2)
    sweep before finally spins ended up in right spin state again.

    To summarize: two complete sweeps (four flips) happened in total.

    NOTE: NMR/EPR toggle switch is getting flaky -> always watch oscilloscope
    and check whether RF is really on.